find the distance between the pair of points (-1,-4) and (-5,-7).
Find the midpoint of the line segment with the given end points: (2,1) and (6,8).
Find the distance between thepair of points (-2,-6) and (7,-3).
Write the standaed form of the quation of the circle with the given center (-2,-4) and radius 6.

Answer by sachi(548)

find the distance between the pair of points (-1,-4) and (-5,-7).
ANS:- D=SQRT{[(-1)-(-5)]^2+[(-4)-(-7)]^2}=SQRT{9+9}=3SQRT(2)
Find the midpoint of the line segment with the given end points: (2,1) and (6,8).
ANS:-M=[(6-2)/2,(8-1)/2]=(2,3.5)
Find the distance between thepair of points (-2,-6) and (7,-3).
ANS:- D=SQRT{[(-2)-(7)]^2+[(-6)-(-3)]^2}=SQRT{81+9}=3SQRT(10)
Write the standaed form of the quation of the circle with the given center
(-2,-4) and radius 6.
ANS:-[x-(-2)]^2+[y-(-4)]^=6^2
or[x+2]^2+[y+4]^2=36
can be simplified to
x^2+4x+y^2+8y=16 ans